OpenStudy (anonymous):

Solve: 2(x + 1) = 2x + 5 options below Any Real Number No Solution 0 3

OpenStudy (anonymous):

2(x + 1) = 2x + 5 2=5 which is a contradiction therefore ,no solution

OpenStudy (anonymous):

2x+2=2x+5 it seems x cancels out so...